### v1

@@ 2. Problem Formulation and Design Ideas
Figure 1. PoD: a consensus protocol for collaborative intelligence.
[…]
PoD consensus
@@ 3.1. Block Structure
  • $ID\_Sender$: unique identity of the sender;
  • $data\_Summary$: a summary that summarizes the characteristics of the data; In our work, we use the Gaussian Mixture Model (i.e., GMM (Bond et al., 2001)) to fit the private datasets of users and $data\_Summary$ mainly includes mean $\mu$, covariance $\Sigma$ and coefficient $\alpha$.
  • $data\_Signature$: a data falsification verification proving the reliability of the training results.
@@ 3.4. Data Verification
One important task of the voting level is data verification. If a training device violates a rule, the voting layer should detect the violation and know which device violated the rule. Accountability allows us to penalize malfeasant devices. For example, a training device has private data $D$=$\{d_{1},d_{2},\cdots,d_{n}\}$, and claims the Gaussian distributions $G$=$\{g_{1},g_{2},\cdots,g_{q}\}$ for all features $F$=$\{f_{1},f_{2},\cdots,f_{q}\}$ in its data. Our data verification objective is to check whether the training devices have data that match the distributions they claim. Meanwhile, we can access devices’ private data.
Concretely, we need to verify the matching for all features. Then, the system splits $f_{j}$’s value range($[Min(f_{j}),Max(f_{j})]$) into some intervals, and the server calculates each interval’s average value according to the corresponding Gaussian distribution $g_{j}$. If the user has data that matches the Gaussian distribution, then the same interval’s average sampled by the user should be similar to the average calculated by the server. To do this, our data verification consists of two parts, namely Consistency Check and Distribution Matching Check
Figure 2. Structure of data verification with privacy protection.
@@
data verification
@@ 3.4.1. Consistency Check
output : Consistency check result $Pass$-$Flag$
1 $Pass$-$Flag$ $\leftarrow True$;
@@ 6 $Send(CM_{k})\rightarrow Server,Peer$;
[…]
      10if Not $Pass$-$Flag$ then
@@ 11 Break;
12return $Pass$-$Flag$;
@@ 3.4.2. Distribution Matching Check
input : $u$ set of all data $U$; $v$ set of all data $V$; system parameter $\phi$; initial summation $A$.
@@ 5 $range=(max-min)/n_{s}$; /*Range of each slice*/
            7 $n_{s}$ = $count_{i}*r_{s}$; /*Number of samples for user $i$/
            8 $p_{i}=\{s_{j}:n_{j}\}$ = random_samples($G_{ik}$, $g_{r}$,$n_{s}$);
            9 /*$s_{j}$: number of slice; */
            10 /*$n_{j}$: number of samples in slice $s_{j}$;*/
      11$\{r_{1k},r_{2k},...,r_{nk}\}$ = feature_contribution($\{p_{1},p_{2},...,p_{n}\}$);
@@ 3.6. An Example for flow of PoD Algorithm
Figure 3. Example of PoD flow.
[…]
Example of PoD flow with four training nodes.
@@ 5.1.2. Dataset allocation
In our experiments, we aim to verify system performance and system fairness in different data distribution scenarios. We mainly simulate the practical application scenarios from three aspects, i.e., data variation (i.e., static vs. dynamic), data overlap (i.e., data intersection ratio) and data volume (i.e., small vs. large dataset). (i) Data variation: We mainly consider static and dynamic data cases and design data that increase the rate of dynamic situations. (ii) Data overlap: We also consider the intersection of user data. Intersection ratio represents data overlap, and in the previous discussion, the greater the data overlap, the lower the data contribution. (iii) Data volumes: Moreover, the amount of user data can greatly influence consensus results, so we also take data volumes into account during the validation process. Specifically, We use the uneven rate to measure the unbalanced volume distribution of user datasets.
Therefore, we designed two ways for data allocation for each data variation scenario in our experiments, i.e., data volume imbalance and overlap rate difference allocations. (i) data volume imbalance: each user is assigned a different number of training samples $\{D_{i}\}_{i\in[n]}$ with a uniformly random distribution over 1,000 classes, and there is no overlap of any two datasets, i.e., $D_{i}\cap D_{j}=\emptyset,1\leq i<j\leq n$. We use the uneven rate $r_{uneven}=1/n\sqrt{\sum_{i\in[n]}{(count_{i}-avg)^{2}}}$, where $avg=1/n\sum_{i\in[n]}count_{i}$. (ii) Overlap rate difference: each user is assigned the same number of training samples with a uniformly random distribution over 1,000 classes, but there exists an overlap between any two datasets, i.e., $D_{i}\cap D_{j}\neq\emptyset,1\leq i<j\leq n$. We use the overlap rate $r_{overlap}=n_{overlap}/\sum_{i\in[n]}count_{i}$, where $n_{overlap}$ represents the data number of all intersections.
@@ 5.1.3. Neural network structure
For the training models used to perform the image classification tasks, we use the convolutional neural network (CNN) for ImageNet, which contains two 5$\times$5 convolutional layers (each layer is followed with a batch normalization and 2$\times$2 max pooling), a fully connected layer with ReLu activation and a final softmax output layer. The CNN model is mended from (McMahan et al., 2017). Unless otherwise specified, some important hyperparameters in our experiments are set as Table 2.

### v2

@@ 2. Problem Formulation and Design Ideas
Figure 1. PoD: a consensus protocol for collaborative intelligence.:
@@ 3.1. Block Structure
  • ID_SenderID\_Sender: unique identity of the sender;
  • data_Summarydata\_Summary: a summary that summarizes the characteristics of the data; In our work, we use the Gaussian Mixture Model (i.e., GMM (Bond et al., 2001)) to fit the private datasets of users and data_Summarydata\_Summary mainly includes mean $\mu$, covariance $\Sigma$ and coefficient $\alpha$.
  • data_Signaturedata\_Signature: a data falsification verification proving the reliability of the training results.
@@ 3.4. Data Verification
[…]
Figure 2. Structure of data verification with privacy protection.:
@@ 3.4.1. Consistency Check
output : Consistency check result PassPass-FlagFlag
1 PassPass-FlagFlag $\leftarrow True$;
@@ 6 $Send(CM_{k})\rightarrow Server,Peer$;
 7 PassPass-FlagFlag $\mathrel{*}=$ Pair-Consist-Check(Server, Peer, $CM_{k}$);
 8 PassPass-FlagFlag $\mathrel{*}=$ Pair-Consist-Check(User, Server, $X_{k}$);
 9 PassPass-FlagFlag $\mathrel{*}=$ Pair-Consist-Check(User, Peer, $Y_{k}$);
 10if Not PassPass-FlagFlag then
@@ 11 Break;
12return PassPass-FlagFlag;
@@ 3.4.2. Distribution Matching Check
input : $u$ set of all data $U$; $v$ set of all data $V$; system parameter $\phi$; initial summation AA.
@@ 5 $range=(max-min)/n_{s}$; /*Range of each slice*/
    7 $n_{s}$ = $count_{i}*r_{s}$; /*Number of samples for user $i$ */
    8 $p_{i}=\{s_{j}:n_{j}\}$ = random_samples($G_{ik}$, $g_{r}$$n_{s}$);
    9 $s_{j}$: number of slice; */
    10 $n_{j}$: number of samples in slice $s_{j}$;*/
[…]
@@ 3.6. An Example for flow of PoD Algorithm
Figure 3. Example of PoD flow.:
@@ 5.1.2. Dataset allocation
In our experiments, we aim to verify system performance and system fairness in different data distribution scenarios. We mainly simulate the practical application scenarios from three aspects, i.e., data variation (i.e., static vs. dynamic), data overlap (i.e., data intersection ratio) and data volume (i.e., small vs. large dataset). (1) Data variation: We mainly consider static and dynamic data cases and design data that increase the rate of dynamic situations. (2) Data overlap: We also consider the intersection of user data. Intersection ratio represents data overlap, and in the previous discussion, the greater the data overlap, the lower the data contribution. (3) Data volumes: Moreover, the amount of user data can greatly influence consensus results, so we also take data volumes into account during the validation process. Specifically, We use the uneven rate to measure the unbalanced volume distribution of user datasets.
Therefore, we designed two ways for data allocation for each data variation scenario in our experiments, i.e., data volume imbalance and overlap rate difference allocations. (1) data volume imbalance: each user is assigned a different number of training samples $\{D_{i}\}_{i\in[n]}$ with a uniformly random distribution over 1,000 classes, and there is no overlap of any two datasets, i.e., $D_{i}\cap D_{j}=\emptyset,1\leq i<j\leq n$. We use the uneven rate $r_{uneven}=1/n\sqrt{\sum_{i\in[n]}{(count_{i}-avg)^{2}}}$, where $avg=1/n\sum_{i\in[n]}count_{i}$. (2) Overlap rate difference: each user is assigned the same number of training samples with a uniformly random distribution over 1,000 classes, but there exists an overlap between any two datasets, i.e., $D_{i}\cap D_{j}\neq\emptyset,1\leq i<j\leq n$. We use the overlap rate $r_{overlap}=n_{overlap}/\sum_{i\in[n]}count_{i}$, where $n_{overlap}$ represents the data number of all intersections.
@@ 5.1.3. Neural network structure
For the training models used to perform the image classification tasks, we use the convolutional neural network (CNN) for ImageNet, which contains two $\times5$ convolutional layers (each layer is followed with a batch normalization and $\times2$ max pooling), a fully connected layer with ReLu activation and a final softmax output layer. The CNN model is mended from (McMahan et al., 2017). Unless otherwise specified, some important hyperparameters in our experiments are set as Table 2.
@@ 7. Conclusions
Acknowledgements.
This research was supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (Project ID: 22-SIS-SMU-048). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.